# When LLM-based Code Generation Meets the Software Development Process

### v1

@@ III-B Communications Among Agents
One of the most important aspects of LLM agents is how the agents communicate. A recent survey paper [40] shows that one common communication pattern is sequential, i.e., ordered, where one agent communicates to the next in a fixed order. Another pattern is disordered, where multiple agents participate in the conversation. Each agent gets the context separately and outputs the response in a shared buffer. Then, the responses can be summarized and used in the next decision-making process. Based on the software process models and the two aforementioned communication patterns, we implement three interaction models for the agents: , , and .
@@
We follow the intuition behind the Waterfall model and implement as an ordered communication among the agents. Given a programming problem, the problem goes through the requirement analysis, design, implementation, and testing. One thing to note is that the test result from our generated tests is redirected to the developer agent in our implementation of so developers can fix and improve the code.
@@ III-B4 Self-Refinement
We implement self-refinement [23], which tries to refine the LLM-generated result through iterative feedback, to further improve the generated artifacts from every development activity. In all three variations of LCG, we assign other agents to review the generated artifacts for every development activity and provide improvement suggestions. We assign the agents from both the downstream activity and the tester to examine the generated artifacts and provide suggestions. The suggestions are then considered for the re-generation of the artifacts. We include the tester in every development activity to emulate the DevTestOps practice [36], where testers are involved in all development activities and provide feedback on the quality aspects. For example, once a requirement engineer generates a requirement document, both the architect and tester would read the document and provide suggestions for improvement. Then, the requirement engineer will re-generate the requirement document based on the previously generated document and suggestions. At the development and testing activity, the tester will generate a test failure report if any of the LLM-generated tests fail or if the code cannot be executed (e.g., due to syntax error). The test failure report is then given to the developer for bug fixing. We repeat the process t times to self-refine the generated code. In our implementation, we currently set t=3. If the code still cannot pass the test, we repeat the entire software development process.
@@ III-C Implementation and Experiment Settings
Environment. We use GPT3.5 (version gpt-3.5-turbo-1106) as our underlying LLM due to its popularity and wide usage in code generation research [18, 34, 9]. We leverage the APIs provided by OpenAI to interact with GPT. We send prompts using JSON format [39] and send all the conversation history as part of the prompt [12]. We set the temperature value to 0.8 and explore the effect of the temperature value in RQ3. We implemented LCG using Python 3.9 and version 0.28.1 for the OpenAI API.
Benchmark Datasets. We follow prior studies [51, 18, 50] and evaluate the code generation result using four benchmarks: HumanEval, HumanEval-ET, MBPP (Mostly Basic Python Programming), and MBPP-ET. These benchmarks contain both the programming problems and tests to evaluate the correctness of the code. Given a programming problem, we consider that a generated code snippet is correct if it can pass all the provided tests. HumanEval [6] has 164 programming problems, and MBPP [2] has 427 programming problems and three test cases for each problem. We also use the dataset published by Dong et al. [8], where they use the same problems as HumanEval and MBPP but offer stronger evaluation test cases (around 100 test cases for each problem, called HumanEval-ET and MBPP-ET). All these benchmarks use Python as the programming language. Each programming problem contains the INPUT pairs method signature, method description, invoke examples and expect the code as OUTPUT.
Evaluation Metric. To evaluate the quality of the generated code, we use the Pass@K metric [9, 6]. Pass@K evaluates the first K generated code’s functional accuracy (i.e., whether the generated code can pass all the test cases). In this work, we set K=1 to evaluate if the first generated code can pass all the provided test cases. A Pass@1 of 100 means 100% of the generated code can pass all the tests in the first attempt.
@@ IV Results
In this section, we present the research questions aimed at evaluating LCG. For each RQ, we discuss the motivation, the approach we take to answering it, and the results and discussion of the findings.
RQ1: What is the code generation accuracy of LCG?
@@
Approach. As a baseline for comparison, we directly give the programming problems to ChatGPT (which we refer to as GPT). Although prior works [15, 17, 25, 18] shows that few-shot learning can improve the results from LLMs, there can be biases on how the few-shot samples are selected [42]. Hence, we use zero-shot learning in our experiment. To control for randomness in the experiment, we ensure all these experiments use the same temperature value (t=0.8) and the same model version (gpt-3.5-turbo-1106). Finally, we repeat each LCG five times and report the average Pass@1 and standard deviation across the runs. We also conduct a student’s t-test to study if LCG’s results are statistically significantly different from GPT.
@@
Result. shows a consistent improvement over GPT, achieving 5.2% to 31.5% improvement in Pass@1. Table II shows the Pass@1 accuracy of LCG across different process models on the benchmark datasets studied. As shown in the Table II, for HumanEval and HumanEval-ET, all of the studied process models have 7.9% to even 31.5% improvement in Pass@1 compared to GPT, and the improvements are all statistically significant (p 0.05). For MBPP and MBPP-ET, also has statistically significant improvements of 5.2% to 6.5%, even though we see a slight decrease when adopting and to MBPP and MBPP-ET.
Despite slight variations in code generation responses form LLM across executions, we find stable standard deviations of Pass@1, ranging from 0.6% to 3.7% across all process models and benchmarks. In particular, has the lowest standard deviation (0.6% to 1.9%, an average of 1.2%), while GPT has the highest standard deviation (0.5% to 3.7%, an average of 2%). Following , has the second highest standard deviation, with is ranking third. In conclusion, although the models generally have consistent Pass@1 across runs, consistently produces the most stable results.
@@
There are potential issues in the tests provided by the benchmarks, which may hinder the Pass@1 of LCG. Table III provides a breakdown of failure types from the Python Interpreter [28] across various process models and benchmarks. While we repeat our experiment 5 times, the standard deviation across runs is low; hence, we represent test failure from only one of the runs. Aligned with the findings from Table II, has the lowest AssertionError compared to other models. We also notice that SyntaxError is more evident in GPT across all benchmarks, as expected due to the absence of a code review and testing process. However, there are still higher test failures in and caused by increased occurrences of ValueError, TypeError, and NameError, for MBPP and MBPP-ET as seen in Table II.
Upon manual investigation of test failures, we discover that and introduce various validations and enforce programming naming conventions in the generated code, which may help improve code quality but cause tests to fail. For example, Listing 1 depicts the provided tests and the generated code/test for MBPP-582, of which the objective is to “Write a function to check if a dictionary is empty or not”. While GPT passed the provided tests, and failed. This failure is because our generated code contains validation to check that the input should be of type dict. However, the MBPP-582 provided test uses an input of the type set, which fails the validation, causing a TypeError exception. Moreover, and enforce the common naming convention format and more meaningful function name (e.g., my_dict v.s. is_dict_empty), both of which causes NameError exception due to missing declaration, causing test failure. More interestingly, we find that such code standardization may also be misled by the requirement provided by the benchmark itself. For example, the MBPP-582 requirement specifies expected input as dict, yet provides a set type as the input to the test. The LLM code generation indeed captures this correct requirement by validating that input must be of type dict. Such inconsistency in the benchmark may reduce the Pass@1.
@@
We conduct a pilot study on the impact of Wrong Function Name and Strict Input Validation by manually fixing these issues in the generated code. We find that, after fixing these issues, most generated code can pass all the tests provided by the benchmarks, indicating correct functional behaviors. Fixing the issues largely improve LCG’s Pass@1: 16 to 28 improvement in HumanEval, 29 to 35 in HumanEval-ET, 15 to 21 in MBPP, and 28 to 42 in MBPP-ET. Namely, LCG’s Pass@1 can achieve over 90 to 95 across all benchmarks. Even though we also observe improvement in GPT’s Pass@1, the three LCG models had greater improvement. On average, the three models have a Pass@1 that is 14%, 8%, 1.4%, and 5% better than GPT in HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively. The aggregated results underscore the deficiencies in the benchmarks, suggesting that the current Pass@1 score of LCG (Table II) could represent a lower bound. These preliminary findings highlight the potential of LCG and suggest that future research should improve the benchmarks by incorporating input checking or consistent naming convention into the tests and subsequently re-evaluate existing code generation techniques.
@@
Approach. To study the impact of each development activity on code quality, we remove each activity separately and re-execute LCG. For example, we first remove the requirement activity in and execute . Then, we add the requirement activity back and remove the design activity. We repeat the same process for every development activity. Note that we cannot remove the coding activity since our goal is code generation. Hence, we removed code review at the end of the coding activity.
Code quality considers numerous facets beyond mere functional correctness [46, 44]. Other factors such as code smells, maintainability, and readability are also related to code quality. Hence, to gain a comprehensive understanding of how code quality changes, in addition to Pass@1, we 1) apply a static code analyzer to detect code smells in the generated code and 2) study code reliability by analyzing the exception handling code. To study code smell, design, and readability, we apply Pylint 3.0.4 [35, 7] (a Python static code analyzer) on the generated code. Pylint classifies the detected code smells into different categories such as error, warning, convention, and refactor.
@@
Result. Testing has the largest impact on the functional correctness of the code, while other development activities only have small impacts. Removing testing causes Pass@1 to decrease by 17.0% to 56.1%. Table IV presents changes in Pass@1 and the densities of code smell and handled exceptions. We show the results for HumanEval and MBPP because they share the same programming problems and generated code with the other two benchmarks. Among all development activities, testing has the largest impact on Pass@1, where removing testing causes a large decrease in Pass@1 (17.0% to 56.1% decrease). The finding implies that LLM’s generated tests are effective in improving the functional correctness of code. In both benchmarks, removing sprint meetings in also causes Pass@1 to drop. However, removing other activities only has a small and inconsistent effect on Pass@1. For example, in HumanEval, removing requirement, design, and code review generally causes Pass@1 to decrease (except for ), but removing these activities improves Pass@1 in MBPP. The findings show that, in HumanEval and MBPP, most development activities do not significantly contribute to the functional correctness of the generated code.
@@
We further study how removing different activities impacts the code smell densities. As shown in Table IV, eliminating test activities significantly boosts error and warning smell densities by an average of 702.2% and 52.2%, respectively. Omitting design raises refactor smell density by an average of 7.1%, and skipping code review leads to a 14.2% average increase in warning density. However, removing other development activities shows either a small or inconsistent impact. In short, the findings show that adding design, testing, and code review can help reduce the density of code smell in the generated code.
@@
Having design and code review activities significantly improves reliability by increasing the density of handled exceptions, while other development activities only have small or no impacts. Removing design and code review activities separately causes the handled exception density to decrease by 20.0% to 43.2% and 14.5% to 27.3%, respectively. Namely, design and code review activities add exception handling in the generated code, which may help improve reliability. Removing other development activities shows a mixed relationship with the density of the handled exception. For example, removing testing in causes an increase of handled exception density by 25.8% in MBPP (i.e., testing removes exception handling code) while causing a decrease of 30.3% in HumanEval (i.e., testing adds exception handling code). While the effect of each development may be related to the nature of the benchmarks, our findings show that, in both benchmarks, adding design and code review activities can help improve code reliability by handling more exceptions in the generated code. LCG shows consistent improvement over GPT in the quality of the generated code: decreasing the density of error/warning/convention/refactor code smells (6.7% to 96.0%) while significantly increasing handled exception density. The code generated by GPT has higher error/warning/convention/refactor code smell densities than that of , , and . This finding shows all three models improve the quality of the generated code to different degrees. Specifically, compared to GPT, LCG decreases the error code smell density by 81.8% to 96.0%, warning density by 15.0% to 57.9%, convention density by 15.4% to 60.2%, and refactor density by 6.7% to 30.0%. Meanwhile, GPT has fewer handled exceptions than LCG. As Table IV shows, in both HumanEval and MBPP, GPT has almost zero handled exception, while generates the most handled-exception (0.37 and 0.67 handled exceptions per every 10 LOC in the two benchmarks), ranks second (0.33 and 0.62), and then (0.15 and 0.44). In conclusion, these findings show that LCG improves the quality of the generated code by reducing code smells while adding more exception-handling code.
RQ3: How stable is the LCG generated code?
@@
Motivation. In LLM, the stability of generated responses can be influenced by several parameters: 1) temperature, affecting the randomness in the generated responses, and 2) model versions, which may introduce variability due to changes in optimization and fine-tuning [5]. Understanding and improving the stability of LLMs is crucial for enhancing their trustworthiness, thereby facilitating their adoption in practice. Therefore, in this research question (RQ), we investigate the stability of our LCG in Pass@1 across four benchmarks, considering various temperature values and model versions.
@@
Approach. We evaluate the Pass@1 of LCG across four versions of GPT3.5 – gpt-3.5-turbo-0301, gpt-3.5-turbo-0613, gpt-3.5-turbo-1106, and gpt-3.5-turbo-0125. The latest version is gpt-3.5-turbo-0125 (published in January 2024) and the earlier version is gpt-3.5-turbo-0301 (published in March 2023). To avoid the effect of the model version when we vary the temperature, we use the same model version (gpt-3.5-turbo-1106, the version that we used in prior RQs) to study the effect of temperature values. We set the temperature to 0.2, 0.4, 0.6, and 0.8 in our experiment. We execute GPT and the three variants of LCG three times under each configuration and report the average Pass@1.
@@
Result. GPT has extremely low Pass@1 in some versions of GPT3.5, while LCG has stable results across all versions. LCG may help ensure the stability of the generated code even when the underlying LLM regresses. Figure 2 shows the Pass@1 for GPT and the three LCG across GPT3.5 versions. In earlier versions of GPT3.5 (0301 and 0613), GPT has very low Pass@1 on all benchmarks (e.g., 20 to 30 in HumanEval and HumanEval-ET). In 0301, MBPP’s Pass@1 is even lower with a value around 5. The findings show that model version may have a significant impact on the generated code. However, we see that, after adopting our agent-based techniques, all three variants of LCG achieve similar Pass@1 across GPT3.5 versions. The results indicate that LCG can generate similar-quality code even if we have an underperformed baseline model.
@@
All techniques have a relatively similar Pass@1 when the temperature value changes. Figure 3 shows the Pass@1 for all the techniques when the temperature value changes. There is a slight downward trend for GPT when increases, but the changes are not significant (Pass@1 is decreased by 2 to 5). For LCG, and especially , we see similar Pass@1 regardless of the temperature value. Although we see a slight increase in the Pass@1 of when =0.4 (2 to 5 higher compared to when =0.8 across the benchmarks), the difference is small, and the Pass@1 is almost the same when is either the lowest (0.2) or largest value (0.8). In short, although temperature values may have an impact on the generated code, the effect is relatively small, especially for LCG.
V Threats to Validity

### v2

@@ III-B Communications Among Agents
One of the most important aspects of LLM agents is how the agents communicate. A recent survey paper [42] shows that one common communication pattern is sequential, i.e., ordered, where one agent communicates to the next in a fixed order. Another pattern is disordered, where multiple agents participate in the conversation. Each agent gets the context separately and outputs the response in a shared buffer. Then, the responses can be summarized and used in the next decision-making process. Based on the software process models and the two aforementioned communication patterns, we implement three interaction models for the agents: , , and (Figure 1). The details of our multi-agent communication history are available online [2].
@@
follows the Waterfall model and implements an ordered communication among the agents. Given a programming problem, the problem goes through the requirement analysis, design, implementation, and testing. One thing to note is that the test result from our generated tests is redirected to the developer agent in our implementation of so developers can fix and improve the code.
@@ III-B4 Self-Refinement
We implement self-refinement [25], which tries to refine the LLM-generated result through iterative feedback, to further improve the generated artifacts from every development activity. In all three variations of FlowGen, we assign other agents to review the generated artifacts for every development activity and provide improvement suggestions. We assign the agents from both the downstream activity and the tester to examine the generated artifacts and provide suggestions. The suggestions are then considered for the re-generation of the artifacts. We include the tester in every development activity to emulate the DevTestOps practice [38], where testers are involved in all development activities and provide feedback on the quality aspects. For example, once a requirement engineer generates a requirement document, both the architect and tester would read the document and provide suggestions for improvement. Then, the requirement engineer will re-generate the requirement document based on the previously generated document and suggestions. At the development and testing activity, the tester will generate a test failure report if any of the LLM-generated tests fail or if the code cannot be executed (e.g., due to syntax error). The test failure report is then given to the developer for bug fixing. We repeat the process t times to self-refine the generated code. In our implementation, we currently set t=3. If the code still cannot pass the test, we repeat the entire software development process.
@@ III-C Implementation and Experiment Settings
Environment. We use GPT3.5 (version gpt-3.5-turbo-1106) as our underlying LLM due to its popularity and wide usage in code generation research [20, 36, 11]. We leverage OpenAI’s APIs (version 0.28.1) to interact with GPT. We send prompts using JSON format [41] and send all the conversation history as part of the prompt [14]. We set the temperature value to 0.8 and explore the effect of the temperature value in RQ3. We implemented FlowGen using Python 3.9.
Benchmark Datasets. We follow prior studies [53, 20, 52] and evaluate the code generation result using four benchmarks: HumanEval, HumanEval-ET, MBPP (Mostly Basic Python Programming), and MBPP-ET. These benchmarks contain both the programming problems and tests for evaluation. Given a programming problem, we consider that a generated code snippet is correct if it can pass all the provided tests. HumanEval [8] has 164 programming problems, and MBPP [3] has 427 programming problems (we use the sanitized version released by the original authors) and three test cases for each problem. We also use the dataset published by Dong et al. [10], where they use the same problems as HumanEval and MBPP but offer stronger evaluation test cases (around 100 test cases for each problem, called HumanEval-ET and MBPP-ET). All these benchmarks use Python as the programming language. Each programming problem contains the INPUT pairs method signature, method description, invoke examples and expect the code as OUTPUT.
Evaluation Metric. To evaluate the quality of the generated code, we use the Pass@K metric [11, 8]. Pass@K evaluates the first K generated code’s functional accuracy (i.e., whether the generated code can pass all the test cases). In this work, we set K=1 to evaluate if the first generated code can pass all the provided test cases. A Pass@1 of 100 means 100% of the generated code can pass all the tests in the first attempt. We use Pass@1 because it is a stricter criterion, reflecting situations where developers do not have the groundtruth for automatically evaluating multiple attempts.
@@ IV Results
We evaluate FlowGen with four research questions (RQs).
RQ1: What is the code generation accuracy of FlowGen?
@@
Approach. As a baseline for comparison, we directly give the programming problems to ChatGPT (which we refer to as RawGPT). Although prior works [17, 19, 27, 20] show that few-shot learning can improve the results from LLMs, they can be biased on how the few-shot samples are selected [44].. Hence, we use zero-shot learning in our experiment. To control for randomness in the experiment, we ensure all these experiments use the same temperature value (t=0.8) and the same model version (gpt-3.5-turbo-1106). Finally, we repeat each FlowGen five times and report the average Pass@1 and standard deviation across the runs. We also conduct a student’s t-test to study if FlowGen’s results are statistically significantly different from RawGPT.
@@
Result. shows a consistent improvement over RawGPT, achieving 5.2% to 31.5% improvement in Pass@1. Table II shows the Pass@1 accuracy of FlowGen across different process models on the benchmark datasets studied. As shown in the Table II, for HumanEval and HumanEval-ET, all of the studied process models have 7.9% to even 31.5% improvement in Pass@1 compared to RawGPT, and the improvements are all statistically significant (p 0.05). For MBPP and MBPP-ET, also has statistically significant improvements of 5.2% to 6.5%, even though we see a slight decrease when adopting and to MBPP and MBPP-ET.
Despite slight variations in code generation responses form LLM across executions, we find stable standard deviations of Pass@1, ranging from 0.6% to 3.7% across all process models and benchmarks. In particular, has the lowest standard deviation (0.6% to 1.9%, an average of 1.2%), while RawGPT has the highest standard deviation (0.5% to 3.7%, an average of 2%). Following , has the second highest standard deviation, with is ranking third. In conclusion, although the models generally have consistent Pass@1 across runs, consistently produces the most stable results.
@@
There are potential issues in the tests provided by the benchmarks, which may hinder the Pass@1 of FlowGen. Table III provides a breakdown of failure types from the Python Interpreter [30] across various process models and benchmarks. For example, IndexError happens when the generated code does not handle an out-of-bound index, causing an exception to be thrown. While we repeat our experiment 5 times, the standard deviation across runs is low; hence, we represent test failure from only one of the runs. Aligned with the findings from Table II, has the lowest AssertionError compared to other models (i.e., higher pass rate). We also notice that SyntaxError is more evident in RawGPT , as expected due to the absence of a code review and testing process. However, there are still higher test failures in and caused by increased occurrences of ValueError, TypeError, IndexError and NameError, for MBPP and MBPP-ET as seen in Table III. Upon manual investigation of prevalent test failures types, we discover that and introduce various validations and enforce programming naming conventions in the generated code, which may help improve code quality but cause tests to fail. For example, Listing 1 depicts the provided tests and the generated code for MBPP-582, of which the objective is to “Write a function to check if a dictionary is empty or not”. While RawGPT passed the provided tests, and failed. This failure is because the generated code contains strict input validation to check that the input should be of type dict. However, the MBPP-582 provided test uses an input of the type set, which fails the validation, causing a TypeError exception. Moreover, and enforce the common naming convention format and more meaningful function name (e.g., my_dict v.s. is_dict_empty), both of which causes NameError exception due to wrong function declaration, causing test failure. More interestingly, we find that such code standardization may also be misled by the requirement provided by the benchmark itself. For example, the MBPP-582 requirement specifies expected input as dict, yet provides a set type as the input to the test. The LLM code generation indeed captures this correct requirement by validating that input must be of type dict. Such inconsistency in the benchmark may reduce the Pass@1.
@@
Fixing the issues largely improve FlowGen’s Pass@1: 16 to 28 improvement in HumanEval, 29 to 35 in HumanEval-ET, 15 to 21 in MBPP, and 28 to 42 in MBPP-ET. Namely, FlowGen’s Pass@1 can achieve over 90 to 95 across all benchmarks. Even though we also observe improvement in RawGPT’s Pass@1, the three FlowGen models had greater improvement. On average, the three models have a Pass@1 that is 14%, 8%, 1.4%, and 5% better than RawGPT in HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively.
The results underscore the deficiencies in the benchmarks, suggesting that the current Pass@1 score of FlowGen could represent a lower bound. These preliminary findings highlight the potential of FlowGen and suggest that future research should improve the benchmarks by incorporating input checking or consistent naming convention into the tests and subsequently re-evaluate existing code generation techniques.
@@
Approach. To study the impact of each development activity on code quality, we remove each activity separately and re-execute FlowGen. For example, we first remove the requirement activity in and execute . Then, we add the requirement activity back and remove the design activity. We repeat the same process for every development activity. Note that we cannot remove the coding activity since our goal is code generation. Hence, we removed code review at the end of the coding activity.
Code quality considers numerous facets beyond mere functional correctness [48, 46]. Other factors, such as code smells, maintainability, and readability, are also related to code quality. Hence, to gain a comprehensive understanding of how code quality changes, we 1) apply a static code analyzer to detect code smells in the generated code and 2) study code reliability by analyzing the exception handling code. To study code smell, design, and readability, we apply Pylint 3.0.4 [37, 9] (a Python static code analyzer) on the generated code. Pylint classifies the detected code smells into different categories such as error, warning, convention, and refactor.
@@
Result. Testing has the largest impact on the functional correctness of the code, while other development activities only have small impacts. Removing testing causes Pass@1 to decrease by 17.0% to 56.1%. Table IV presents changes in Pass@1 and the densities of code smell and handled exceptions. We show the results for HumanEval and MBPP because they share the same programming problems and generated code with the other two benchmarks. Among all development activities, testing has the largest impact on Pass@1, where removing testing causes a large decrease in Pass@1 (17.0% to 56.1% decrease). The finding implies that LLM’s generated tests are effective in improving the functional correctness of code. In both benchmarks, removing sprint meetings in also causes Pass@1 to drop. However, removing other activities only has a small and inconsistent effect on Pass@1. For example, in HumanEval, removing requirement, design, and code review generally causes Pass@1 to decrease (except for ), but removing these activities improves Pass@1 in MBPP. In other words, most development activities do not significantly contribute to the functional correctness of the generated code.
@@
As shown in Table IV, eliminating test activities significantly boosts error and warning smell densities by an average of 702.2% and 52.2%, respectively. Omitting design raises refactor smell density by an average of 7.1%, and skipping code review leads to a 14.2% average increase in warning density. However, removing other development activities shows either a small or inconsistent impact. We also find some differences in the artifacts generated by different roles. For example, although both roles generate documents, requirement engineers specify the acceptance criteria, while architects address time/space complexity. In short, the findings show that adding design, testing, and code review can help reduce the density of code smell in the generated code.
@@
Having design and code review activities significantly improves reliability by increasing the density of handled exceptions, while other development activities only have small or no impacts. Removing design and code review activities separately causes the handled exception density to decrease from 20.0% to 43.2% and 14.5% to 27.3%, respectively. Namely, these two activities add exception handling in the generated code, which may help improve reliability. Removing other activities shows a mixed relationship with the density of the handled exception. For example, removing testing in causes an increase of handled exception density by 25.8% in MBPP (i.e., testing removes exception handling code) while causing a decrease of 30.3% in HumanEval (i.e., testing adds exception handling code). While the effect of each development may be related to the nature of the benchmarks, our findings show that, in both benchmarks, adding design and code review activities can help improve code reliability by handling more exceptions in the generated code.
FlowGen shows consistent improvement over RawGPT in the quality of the generated code: decreasing the density of error/warning/convention/refactor code smells (6.7% to 96.0%) while significantly increasing handled exception density. The code generated by RawGPT has higher error/warning/convention/refactor code smell densities than that of , , and . This finding shows all three models improve the quality of the generated code to different degrees. Specifically, compared to RawGPT, FlowGen decreases the error code smell density by 81.8% to 96.0%, warning density by 15.0% to 57.9%, convention density by 15.4% to 60.2%, and refactor density by 6.7% to 30.0%. Meanwhile, RawGPT has fewer handled exceptions than FlowGen. As Table IV shows, in both HumanEval and MBPP, RawGPT has almost zero handled exception, while generates the most handled-exception (0.37 and 0.67 handled exceptions per every 10 LOC in the two benchmarks), ranks second (0.33 and 0.62), and then (0.15 and 0.44). In short, FlowGen improves the quality of the generated code by reducing code smells while adding more exception-handling code.
RQ3: How stable is the FlowGen generated code?
@@
Motivation. In LLM, the stability of generated responses can be influenced by several parameters: 1) temperature, affecting the randomness in the generated responses, and 2) model versions, which may introduce variability due to changes in optimization and fine-tuning [7]. Understanding and improving the stability of LLMs is crucial for enhancing their trustworthiness, thereby facilitating their adoption in practice. Therefore, in this research question (RQ), we investigate the stability of our FlowGen in Pass@1 across four benchmarks, considering various temperature values and model versions.
@@
Approach. We evaluate the Pass@1 of FlowGen across four versions of GPT3.5: turbo-0301, turbo-0613, turbo-1106, and turbo-0125. The latest version is turbo-0125 (published in January 2024), and the earlier version is turbo-0301 (published in March 2023). To avoid the effect of the model version when we vary the temperature, we use the same model version (turbo-1106, the version that we used in prior RQs) to study the effect of temperature values. We set the temperature to 0.2, 0.4, 0.6, and 0.8 in our experiment. We execute RawGPT and the three variants of FlowGen three times under each configuration and report the average Pass@1.
@@
Result. RawGPT has extremely low Pass@1 in some versions of GPT3.5, while FlowGen has stable results across all versions. FlowGen may help ensure the stability of the generated code even when the underlying LLM regresses. Figure 2 shows the Pass@1 for RawGPT and the three FlowGen across GPT3.5 versions. In earlier versions of GPT3.5 (0301 and 0613), RawGPT has very low Pass@1 on all benchmarks (e.g., 20 to 30 in HumanEval and HumanEval-ET). In 0301, MBPP’s Pass@1 is even lower with a value around 5. The findings show that model version may have a significant impact on the generated code. However, we see that, after adopting our agent-based techniques, all three variants of FlowGen achieve similar Pass@1 across GPT3.5 versions. The results indicate that FlowGen can generate similar-quality code even if we have an underperformed baseline model.
@@
All techniques have a relatively similar Pass@1 when the temperature value changes. Figure 3 shows the Pass@1 for all the techniques when the temperature value changes. There is a slight downward trend for RawGPT when increases, but the changes are not significant (Pass@1 is decreased by 2 to 5). For FlowGen, and especially , we see similar Pass@1 regardless of the temperature value. Although we see a slight increase in the Pass@1 of when =0.4 (2 to 5 higher compared to when =0.8 across the benchmarks), the difference is small, and the Pass@1 is almost the same when is either the lowest (0.2) or largest value (0.8). In short, although temperature values may have an impact on the generated code, the effect is relatively small for FlowGen.
RQ4: How does FlowGen compare with other code generation techniques?
Motivation. FlowGen is designed to organize agents to emulate process models and can be combined with other code generation techniques. However, it is crucial to evaluate its performance relative to these techniques to assess the effectiveness in emulating software process models for code generation. While many code generation results use the same benchmarks, our evaluation results cannot be directly compared with other agent-based code generation works [14, 15, 16, 5, 36] due to missing information on model versions, temperature values, post-processing steps, specific prompts, or the selection of few-shot samples. Therefore, in this RQ, we compare with other LLM-based baselines under the same environment settings. Moreover, we evaluate an integrated version of to showcase how existing prompting techniques can be combined with FlowGen.
Approach. We compare against two state-of-the-art techniques: CodeT [5] and Reflexion [36]. CodeT employs self-generated tests to evaluate the quality of generated code, which is similar to the testing phase of FlowGen. Reflexion is an agent-based technique that achieves state-of-the-arts Pass@1 on the benchmarks. We apply these two techniques using their released code, replacing the LLM version and temperature settings with those used by , and repeat the experiment five times.
Result. While and CodeT achieve similar results, with having a higher Pass@1 in MBPP, they both have higher Pass@1 than Reflexion on all benchmarks (statistically significant). Table V shows the Pass@1 of the techniques. has a statistically significantly higher Pass@1 than CodeT in MBPP and similar Pass@1 in other benchmarks (no statistically significant difference). Both techniques achieve higher Pass@1 than Reflexion (statistically significant). Reflexion’s MBPP results are even worse than RawGPT. This observation aligns with the original study [36], where Shinn et al. reported similar performance degradation.
Integrating CodeT to further brings statistically significant improvements of Pass@1 by up to 5%. CodeT is a general technique where it repeats the code generation and selects the code that passes the most self-generated tests. Hence, as a pilot study, we made the developer agent repeats the implementation activity multiple times to produce several versions of the code (i.e., ). The developer agent then generates multiple test assertions to identify the version with the highest pass rate. The selected code is subsequently submitted to the next stage: the testing activity. Our findings indicate that outperforms and CodeT, achieving an average Pass@1 score of 83.8, 58.3, 79.3, and 67.7 on HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively. This provides statistically significant improvement over both and CodeT. Our finding highlights the potential of FlowGen in boosting the performance of other code generation techniques (and vice versa). Future studies can refine FlowGen to incorporate enhancement to each activity for further improvement. To support these efforts, we have made our code publicly available [2] to facilitate further adoption and allow researchers to experiment with different software process models.
| HumanEval | HumanEval-ET | MBPP | MBPP-ET | |
| Reflexion | 71.31.5 | 55.72.8 | 71.70.8 | 52.00.7 |
| CodeT | 75.70.4 | 66.90.4 | 79.91.3 | 56.70.9 |
| 75.21.1 | 65.51.9 | 82.50.6 | 56.71.4 | |
| 79.31.6 | 67.71.1 | 83.80.6 | 58.71.3 |
V Discussion & Future Works
Role of Human Developers in FlowGen. Although software process models were originally designed for human-centric development rather than for LLMs, our empirical findings suggest that certain elements of these processes can contribute to better code quality. Every activity in the process model also has different impacts on the generated artifacts. Future research should examine the incorporation of human developers into various phases of the code generation process. Specifically, humans can play critical roles in the following stages: (1) Pre-Execution of FlowGen: different process models exhibit varying quality (e.g., smell and accuracy). Humans are instrumental in selecting the most appropriate model for a given task. Humans are also essential in providing initial requirements and design specifications. (2) During FlowGen Execution: Humans can oversee review meetings and assist in reviewing/improving generated artifacts. For example, humans can validate the generated requirements with product managers, or verify the quality of the generated code by manual inspection and debugging. The improvement in each activity can also impact the subsequent activities and, hence, affect the final artifacts. (3) Post-Execution of FlowGen: Following the code-generation phase, humans can either accept the generated artifact or request further refinements, offering additional requirements as needed to better meet project goals.
Quality of Code Generation Benchmarks We manually validated all coding problems that failed the tests. We found that the majority of quality issues within the benchmark were in MBPP and MBPP-ET (e.g., bad naming convention or inconsistent test definition). These issues may contribute to reduced Pass@1 scores due to factors beyond logic in the code. It is also important to acknowledge that other benchmarks might present unique challenges that could similarly affect Pass@1 evaluations. Hence, a crucial research direction is to conduct a thorough evaluation of benchmarks for more diverse and accurate evaluations of code generation approaches.
VI Threats to Validity